OpenStudy (anonymous):

What's the equation of the line passing through the point (6, -2)and parallel to y = x + 7.

OpenStudy (aroub):

y=mx+b , m=1 , (6,-2) -2=1(6)+b -2=6+b -8=b =>y=x-8